OpenStudy (anonymous):

PLEASE HELP i AM LOST AT THIS Find the vertex of parabola f(x)=x^2-10-3

OpenStudy (amistre64):

completing the square might be useful; or if you know the vertex part of the quadratic formula

OpenStudy (amistre64):

given: ax^2 + bx + c the x value of the vertex has the form: -b/(2a)

OpenStudy (amistre64):

and i believe the y part would amount to: c - (b/2a)^2

OpenStudy (anonymous):

the problem is x^2-10x-3

OpenStudy (amistre64):

can you define the abc parts?

OpenStudy (anonymous):

yes this is what I have so far on this problem x=- -10/2*1 right

OpenStudy (amistre64):

x = --10/2*1 is good to me, simplifies to 5

OpenStudy (amistre64):

lets make this simpler for y, i have a notion that once you find x; then y= c - x^2

OpenStudy (amistre64):

so, y= -3 - (5)^2

OpenStudy (anonymous):

so it would not be (5)^2-10*5-3 then that is where I get lost

OpenStudy (amistre64):

it would be, but thats a rather brute way to go about it; it gets the same results tho if i take my idea from the method of completing the square: ax^2 + bx + c x^2 + b/a x + c/a x^2 + b/a x +(b/2a)^2 -(b/2a)^2 + c/a (x^2 + b/a x +(b/2a)^2) -(b/2a)^2 + c/a (x - -(b/2a))^2 + ( c/a - (b/2a)^2 ) ^^ ^^^^^^^^^^^^ X vertex Y vertex so, x=-b/2a ; and y= c/a - x^2 .... i was close, but since a=1 it was immaterial

OpenStudy (anonymous):

so the vertex of the fiven parabola is (5,??) what is the last number?

OpenStudy (anonymous):

given sorry

OpenStudy (amistre64):

y= -3 - (5)^2 is what i keep posting :)

OpenStudy (amistre64):

if we go the brute method (5)^2: 25 -10*5:-50 -3: -3 ---------- y : 25-53 im really not sure where the confusion is at